OpenStudy (anonymous):

Find the number of sides of each of the two polygons if the total number of sides of the polygons is 13, and the sum of the number of diagonals of the polygons is 25.

OpenStudy (jack1):

apparently: number of diagonals in a polygon = 1/2 N(N-3) and N = number of sides of 1 polygon so call them a and b a + b =13 sides total 1/2 a(a-3) + 1/2 b(b-3) = 25

OpenStudy (jack1):

so simultaneous equations, rearrange one in terms of the other and sub it into the other equation

OpenStudy (jack1):

so if a + b = 13 a = 13 - b so now u can substitute (13-b) into the 2nd equation, wherever you see an "a".... im not 100% on this tho dude, let me know how u go, yeah?

OpenStudy (jack1):

@klionmaster10 ...?

OpenStudy (anonymous):

Trying to solve.

OpenStudy (anonymous):

Thanks dude! I finally got it! :)

OpenStudy (jack1):

sweet, what were ur answers if u dont mind me asking?

OpenStudy (jack1):

octogon and 5 sided polygon?

OpenStudy (anonymous):

8 and 5

OpenStudy (jack1):

niiiice! glad it worked ;S slaters!